40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Jan 28 17:39:17 2015, Koji, Configuration, IOO, X Trans Table less crazy but not enough yet IMG_1808.JPG
    Reply  Thu Jan 29 17:20:58 2015, manasa, Configuration, IOO, X Trans Table less crazy but not enough yet 
Message ID: 10951     Entry time: Wed Jan 28 17:39:17 2015     Reply to this: 10958
Author: Koji 
Type: Configuration 
Category: IOO 
Subject: X Trans Table less crazy but not enough yet 

The X-end IR Trans path was cleaned up.

I have been investigating the Xarm ASS issue. The Xarm ASS sensors behaved not so straight forward.
I went to the X-end table and found some suspect of clipping and large misalignmnet in the IR trans path.
Facing with the usual chaos of the end table, I decided to clean-up the IR trans path.

The optical layout is now slightly better. But the table is, in general, still dirty with bunch of stray optics,
loose cables and fibers. We need more effort to make the table maintained in a professional manner.


- Removed unnecessary snaking optical path. Now the beam from the 1064/532 separator is divided by a 50-50 BS before the QPD without
any other steering mirrors. This means the spot size on the QPD was changed as well as the alignment. The spot on the QPD was aligned
with the arm aligned with the current (=not modified) ASS. This should be the right procedure as the spot must be centered on the end mirror
with the current ASS.

- After the 50-50 BS there is an HR steering mirror for the Thorlab PD.

- A VIS rejection filter was placed before the 50-50 BS. The reflection from the filter is blocked with a razor blade dump.

Important note to everyone including Steve:
The transmission of the VIS rejection filter at 1064nm is SUPER angular sensitive.
A slight tilt causes significant reduction of 1064nm light. Be careful.

- As we don't need double VIS filter, I removed the filter on the QPD.

- X-End QPD was inspected. There seemed large (+/-10%) gain difference between the segments.
They were corrected so that the values are matched when the beam is only on one segment.
The corrections were applied at C1:SUS-ETMX_QPDx_GAIN (x=1, 2, 3, or 4).


I decided to put "-20dB" filters on C1:SUS-ETMi_QPD_SUM and C1:SUS-ETMi_TRY (i = X or Y)
in order to make their gain to be reasonable (like 0.123 instead 0.000123 which is unreadable).
Jenne's normalization script reads relative values and the current gains instead of the absolute values.
Therefore the script is not affected.

Attachment 1: IMG_1808.JPG  624 kB  Uploaded Wed Jan 28 19:09:31 2015  | Hide | Hide all
IMG_1808.JPG
ELOG V3.1.3-